OpenStudy (anonymous):

The product of (a - b)(a - b) is a perfect square trinomial. Answer choices below :)

OpenStudy (anonymous):

Sometimes Always Never

OpenStudy (anonymous):

its never

OpenStudy (anonymous):

its never

OpenStudy (cwrw238):

(a - b)(a - b) = a^2 -2ab + b^2

OpenStudy (anonymous):

why never? \[(a - b)(a - b) =(a-b)^2\] so of course it is a perfect square it is a square before you start

OpenStudy (anonymous):

Yahoo says always xD Im confused

OpenStudy (anonymous):

it is a square because \(5\times 5=5^2\) and \(x\times x=x^2\) and \((a-b)(a-b)=(a-b)^2\)